OpenStudy (anonymous):

a jet ascends at a 10 degree angle with an airspeed of 550 mph. How fast is the altitude of the jet increasing? If the sun is directly overhead, how fast is the shadow of the jet moving on the ground? calculus question using related rates

OpenStudy (wolf1728):

sine (10°) = x / 550 .17365 = x / 550 x = 550 * .17365 x = 95.5075 mph (rate at which altitude is increasing.

OpenStudy (anonymous):

So i do not have to find dy/dx of anything?

OpenStudy (wolf1728):

Shadow of the Jet cosine 10° = adjacent / 550 .98481 = adj / 550 adj = .98481 * 550 adj = 542 mph = speed of airplane's shadow

OpenStudy (wolf1728):

No, to me it seems you don't need calculus at all.